# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ 3.1 Safe Mean-Field Reinforcement Learning
$h_{C}(\mu_{t+1})$ $\geq 0,$ (4e)
where we explicitly require induced mean-field distributions $\{\mu_{t}\}_{t=1}^{T}$ to reside in the safe set $\zeta$ by restricting the set of admissible policy profiles $\Pi$ to policy profiles that induce safe distributions. To ensure complete safety, we note that the initial mean-field distribution $\mu_{0}$ must be in the safe set $\zeta$ as our learning protocol does not induce $\mu_{0}$ (Algorithm 1) and $s_{0}$ is sampled from $\mu_{0}$.
We make the following assumptions about the environment using the Wasserstein $1$-distance defined as $W_{1}(\mu,\mu^{\prime}):=\inf_{\gamma\in\Gamma(\mu,\mu^{\prime})}\mathbb{E}_{(x,y)\sim\gamma}\|x-y\|_{1}$ where $\Gamma(\mu,\mu^{\prime})$ is the set of all couplings of $\mu$ and $\mu^{\prime}$ (i.e., a joint probability distributions with marginals $\mu$ and $\mu^{\prime}$). We further define the distance between $z=(s,\mu,a)$ and $z^{\prime}=(s^{\prime},\mu^{\prime},a^{\prime})$ as $d(z,z^{\prime}):=\|s-s^{\prime}\|_{2}+\|a-a^{\prime}\|_{2}+W_{1}(\mu,\mu^{\prime})$.
Assumption 1 (Transition dynamics Lipschitz continuity).
The dynamics function $f(\cdot)$ is $L_{f}$-Lipschitz-continuous, i.e., $\|f(z)-f(z^{\prime})\|_{2}\leq L_{f}d(z,z^{\prime})$.
@@
The individual policies present in any admissible policy profile in $\Pi$ are $L_{\pi}$-Lipschitz-continuous, i.e., $\|\pi(s,\mu)-\pi(s^{\prime},\mu^{\prime})\|_{2}\leq L_{\pi}(\|s-s^{\prime}\|_{2}+W_{1}(\mu,\mu^{\prime}))$ for all $\pi\in\bm{\pi}\in\Pi$.
@@
The reward function $r(\cdot)$ is $L_{r}$-Lipschitz-continuous, i.e., $\|r(z)-r(z^{\prime})\|_{2}\leq L_{r}d(z,z^{\prime})$.
These assumptions are considered standard in model-based learning [28, 12, 47, 43] and mild, as individual policies and rewards are typically designed in such a way that they meet these smoothness requirements. For example, we use neural networks with Lipschitz-continuous activations to represent our policies. Therefore the obtained policies are also Lipschitz-continuous when the weights are bounded (see Section D.3).
3.2 Statistical Model and Safety Implications
The representative agent learns about the unknown transition dynamics by interacting with the environment. We take a model-based approach to achieve sample efficiency by sequentially updating and improving the dynamics model estimates based on the previously observed transitions. At the beginning of each episode $n$, the representative agent updates its model based on $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t+1})\}_{t=0}^{T-1}$ and $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$ is the set of observations in episode $i$ for $i=1,...,n-1$, i.e., up until the beginning of episode $n$. We estimate the mean vector $\bm{m}_{n-1}:\mathcal{Z}\rightarrow\mathcal{S}$ and covariance matrix $\bm{\Sigma}_{n-1}:\mathcal{Z}\rightarrow\mathbb{R}^{p\times p}$ from the set of collected trajectories $\cup_{i=1}^{n-1}\mathcal{D}_{i}$, and denote model’s confidence with $\bm{\sigma}_{n-1}(z)=\text{diag}(\bm{\Sigma}_{n-1}(z))$. We assume that the statistical model is calibrated, meaning that at the beginning of every episode, the agent has high probability confidence bounds around unknown dynamics. The following assumptions are consistent with [50, 12, 15, 47, 43] and other literature which aims to exclude extreme functionals from consideration.
@@ 3.2 Statistical Model and Safety Implications
Let $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ be the mean and covariance functions of the statistical model of $f$ conditioned on $n$ observed episodes. For the confidence function $\bm{\sigma}_{n-1}(\cdot)$, there exists a non-decreasing, strictly positive sequence $\{\beta_{n}\}_{n\geq 0}$ such that for $\delta>0$ with probability at least $1-\delta$, we have jointly for all $n\geq 1$ and $z\in\mathcal{Z}$ that $|f(z)-\bm{m}_{n-1}(z)|\leq\beta_{n-1}\bm{\sigma}_{n-1}(z)$ elementwise.
@@
The confidence function $\bm{\sigma}_{n}(\cdot)$ is $L_{\sigma}$-Lipschitz-continuous for all $n\geq 0$, i.e., $\|\bm{\sigma}_{n}(z)-\bm{\sigma}_{n}(z^{\prime})\|_{2}\leq L_{\sigma}d(z,z^{\prime})$.
Since the true transition dynamics model is unknown in Equation 4c and Equation 4d, at the beginning of every episode $n$, the representative agent can only construct the confidence set of dynamics functions $\mathcal{F}_{n-1}$ with $\bm{m}_{n-1}(\cdot)$ and $\bm{\sigma}_{n-1}(\cdot)$ estimated based on the observations up until the end of the previous episode $n-1$, i.e.,
$\mathcal{F}_{n-1}=\Big{\{}\tilde{f}:|\tilde{f}(z)-\bm{m}_{n-1}(z)|\leq\beta_{n-1}\bm{\sigma}_{n-1}(z),\text{ elementwise and }\forall z\in\mathcal{Z}\Big{\}}.$ (5)
The crucial challenge in Safe Mean-Field Reinforcement Learning is that the representative agent can only select dynamics $\tilde{f}\in\mathcal{F}_{n-1}$ at the beginning of the episode $n$ and use it instead of true dynamics $f$ when solving Section 3.1 to find an optimal policy profile $\bm{\pi}_{n}^{*}$. The resulting mean-field distributions $\{\tilde{\mu}_{t}\}_{t=1}^{T}$ are then different from $\{\mu_{t}\}_{t=1}^{T}$ (i.e., the ones that correspond to the true dynamics model), and hence the constraint Equation 3 guarantees only the safety under the dynamics $\tilde{f}$, i.e., $h_{C}(\tilde{\mu}_{t})\geq 0$. In contrast, the original environment constraint $h_{C}(\mu_{t})\geq 0$ might be violated, resulting in unsafe mean-field distributions under the true dynamics $f$.
Next, we demonstrate how to modify the constraint Equation 4e for the optimization problem Section 3.1 when a dynamics function $\tilde{f}$ is used from the confidence set $\mathcal{F}_{n-1}$ such that the mean-field distributions induced by the resulting policy profile do not violate the original constraint under the true dynamics $f$. First, we require the following property for any safety function $h(\cdot)$:
@@
The previous assumption implies that the safety constraint $h_{C}(\cdot)$ exhibits analogous property. The following lemma shows that we can ensure safety under the true dynamics $f$ by having a tighter constraint under any dynamics $\tilde{f}$ selected from $\mathcal{F}_{n-1}$.
@@
Given a fixed policy profile $\bm{\pi}_{n}$, a safety function $h(\cdot)$ satisfying Assumption 6 and a safety threshold $C\in\mathbb{R}$, we have in episode $n$ for all steps $t$:
@@
For arbitrary $\tilde{\mu}_{n,t},\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ we have $|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|=|h(\tilde{\mu}_{n,t})-h(\mu_{n,t})|\leq L_{h}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})\leq L_{h}C_{n,t}$, where the first equality follows from the definition of $h_{C}(\cdot)$, the first inequality follows from Assumption 6 and the second inequality comes from $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$.∎
Crucially, using Lemma 1 we can formulate a safety constraint for the optimization under dynamics $\tilde{f}$ that ensure that the constraint under the true dynamics $f$ is not violated.
@@
For every episode $n$ and every step $t$, $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ implies $h_{C}(\mu_{n,t})\geq 0$ guaranteeing the safety of the original system.
@@
The corollary follows directly from Lemma 1 and the triangle inequality:
$L_{h}C_{n,t}\leq h_{C}(\tilde{\mu}_{n,t})\leq|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|+h_{C}(\mu_{n,t})\leq L_{h}C_{n,t}+h_{C}(\mu_{n,t}),$
where the second inequality follows from the triangle inequality and the third from Lemma 1. The corollary then follows by subtracting the positive constant $L_{h}C_{n,t}$ from both sides. ∎
@@
Then, $C_{n,t}$ for $t=1,\dots,T$ become parameters of the optimization problem (as defined in Section 4) that the representative agent faces at the beginning of episode $n$. However, choosing the appropriate values that comply with the condition $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ is not trivial since $\mu_{n,t}$ depends on the unknown true dynamics of the system. Nevertheless, note that computing $C_{n,0}$ at the initial step $t=0$ is not necessary because the inequality is always guaranteed due to the initialization $\tilde{\mu}_{n,0}=\mu_{n,0}$ for every episode $n$. In Appendix A, we demonstrate how to efficiently upper bound $W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and obtain $C_{n,t}$ using the Lipschitz constants of the system and the statistical model’s epistemic uncertainty. In particular, $C_{n,t}$ approaches zero, and $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ reduces to the constraint Equation 4e as $\bm{\sigma}_{n-1}(\cdot)$ shrinks due to the increasing number of observations available to estimate the true dynamics.
3.3 Examples of Safety Constraints
We can model multiple classes of safety constraints $h_{C}(\cdot)\geq 0$ that naturally appear in real-world applications such as vehicle repositioning, traffic flow, congestion control, and others. We provide concrete examples of safety functions in Appendix B together with proofs that they satisfy Assumption 6.
Entropic safety. Entropic constraints can be used in multi-agent systems to prevent overcrowding by promoting spatial diversity and preventing excessive clustering. Incorporating an entropic term in the decision-making process encourages the central controller to distribute the agents evenly within the state space. This can be particularly useful in applications that include crowd behavior, such as a swarm of drones or a fleet of vehicles. In such scenarios, we define safety by imposing a hard threshold $C>0$ on the entropy of the mean-field distribution $\mu$:
$h_{C}(\mu)=-\int_{\mathcal{S}}\log\mu(s)\mu(ds)-C.$
Distribution similarity. Another way to define safety is by preventing $\mu$ from diverging from a prior distribution $\nu_{0}$. The prior can be based on previous studies, expert opinions or regulatory requirements. We can use a penalty function that quantifies the allowed dissimilarity between the two distributions:
$h_{C}(\mu;\nu_{0})$ $=C-D(\mu,\nu_{0}),$
with $C>0$ and where the distance function between probability measures $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$ depends on the problem at hand.
4 Safe-$\text{M}^{3}$-UCRL
In this section, we introduce a model-based approach for the Safe Mean-Field Reinforcement Learning problem that combines the safety guarantees in Corollary 1 with upper-bound confidence interval optimization. At the beginning of each episode $n$, the representative agent constructs the confidence set of dynamics functions $\mathcal{F}_{n-1}$ (see Equation 5) given the calibrated model and previously observed data and selects a safe optimistic policy profile $\bm{\pi}_{n}^{*}$ to obtain the highest value function within $\mathcal{F}_{n-1}$ while satisfying the safety constraint derived in Corollary 1. In particular, the optimal policy profile $\bm{\pi}^{*}$ from Section 3.1 is approximated at the episode $n$ by:
$\bm{\pi}_{n}^{*}$ $=\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}\max_{\tilde{f}_{n-1}\in\mathcal{F}_{n-1}}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})\Big{\rvert}\tilde{\mu}_{n,0}=\mu_{0}\Bigg{]}$ (6a)
s.t. $\quad\tilde{a}_{n,t}=\pi_{n,t}(\tilde{s}_{n,t},\tilde{\mu}_{n,t})$ (6b)
$\tilde{s}_{n,t+1}=\tilde{f}_{n-1}(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\varepsilon_{n,t}$ (6c)
$\tilde{\mu}_{n,t+1}=U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1})$ (6d)
$h_{C}(\tilde{\mu}_{n,t+1})\geq L_{h}C_{n,t+1}.$ (6e)
Section 4 optimizes over the function space $\mathcal{F}_{n-1}$ which is usually intractable even in bandit settings [16]. Additionally, it must comply with the safety constraint Equation 6e, further complicating the optimization. We utilize the hallucinated control reparametrization and the log-barrier method to alleviate these issues. After the reformulation of the problem, model-free or model-based mean-field optimization algorithms can be applied to find policy profile $\bm{\pi}_{n}^{*}$ at the beginning of episode $n$.
We use the established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) [39, 15, 43] and introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated dynamics model:
Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form Equation 7:
$\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\;\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.$
Thus, the optimization problem in Section 4 can be expressed as an optimization over the set of admissible policies $\Pi$ and auxiliary function $\eta(\cdot)$ (see Section C.2 for further details). Note that $\eta(z)=\eta(s,\mu,\pi(s,\mu))=\eta(s,\mu)$ for a fixed individual policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of dynamics functions $\mathcal{F}_{n-1}$ [15]. Furthermore, Equation 7 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi},\eta(\cdot),\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ using gradient ascent.
We also introduce the safety constraint to the objective using the log-barrier method [55]. This method restricts the domain on which the objective function is defined only to values that satisfy the constraint Equation 6e, hence, turning Section 4 to an unconstrained optimization problem. Combining these two methods yields the following optimization problem:
$\bm{\pi}_{n}^{*}=\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}$ $\max_{\eta(\cdot)\in[-1,1]^{p}}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\lambda\log(h_{C}(\tilde{\mu}_{n,t+1})-L_{h}C_{n,t+1})\Big{\rvert}\tilde{\mu}_{n,0}=\mu_{0}\Bigg{]}$ (8a)
[…]
$\tilde{f}_{n-1}(\tilde{z})$ $=\bm{m}_{n-1}(\tilde{z})+\beta_{n-1}\bm{\Sigma}_{n-1}(\tilde{z})\eta(\tilde{z})$ (8c)
$\tilde{s}_{n,t+1}$ $=\tilde{f}_{n-1}(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\varepsilon_{n,t}$ (8d)
$\tilde{\mu}_{n,t+1}$ $=U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1}),$ (8e)
with $\lambda>0$ being a tuneable hyperparameter used to balance between the reward and the safety constraint. Provided that the set of safe mean-field distributions (including the initial distribution $\mu_{0}$ that we initialize ourselves) is not empty, $\bm{\pi}_{n}^{*}$ is guaranteed not to violate the safety constraint during the policy rollout in episode $n$.
Remark 1.
Note that Section 4 can also be used under the assumption of known true dynamics by setting $\bm{m}_{n-1}(\cdot)=f(\cdot)$, $\bm{\Sigma}_{n-1}(\cdot)=0$ and $L_{h}C_{n,t}=0$, hence, recovering the original constraint $h_{C}(\tilde{\mu}_{n,t})\geq 0$ from Equation 3. In Section 5, we utilize this useful algorithm property to construct the upper bound for the reward obtained with unknown dynamics.
Algorithm 1 Model-based Learning Protocol in Safe-$\text{M}^{3}$-UCRL
@@ 4 Safe-$\text{M}^{3}$-UCRL
3:     Compute $C_{n,t}$ for $t=1,\ldots,T$
4:     Optimize the objective in Section 4 with respect to the policy profile and hallucinated dynamics
5:     Execute the obtained policy profile and collect the trajectories from the representative agent
6:     Update the confidence set of dynamics $\mathcal{F}_{n-1}$ with the collected data to obtain $\mathcal{F}_{n}$ for the next episode

### v2

@@ 3.1. Safe Mean-Field Reinforcement Learning
$h_{C}(\mu_{t{}+{}1})$ ${}\geq{}0,$ (4e)
where we explicitly require induced mean-field distributions $\{\mu_{t}\}_{t=1}^{T}$ to reside in the safe set $\zeta$ by restricting the set of admissible policy profiles $\Pi$ to policy profiles that induce safe distributions. To ensure complete safety, we note that the initial mean-field distribution $\mu_{0}$ must be in the safe set $\zeta$ as our learning protocol does not induce it (see Algorithm 1).
We make the following assumptions about the environment using Wasserstein $1$-distance defined by
$W_{1}(\mu,\mu^{\prime}){}:={}\inf_{\gamma\in\Gamma(\mu,\mu^{\prime})}\mathbb{E}_{(x,y)\sim\gamma}\|x-y\|_{1},$
where $\Gamma(\mu,\mu^{\prime})$ is the set of all couplings of $\mu$ and $\mu^{\prime}$ (i.e., a joint probability distributions with marginals $\mu$ and $\mu^{\prime}$). We further define the distance between $z=(s,\mu,a)$ and $z^{\prime}=(s^{\prime},\mu^{\prime},a^{\prime})$ as
$d(z,z^{\prime}){}:={}\|s-s^{\prime}\|_{2}{}+{}\|a-a^{\prime}\|_{2}{}+{}W_{1}(\mu,\mu^{\prime}).$
Assumption 1 (Transitions Lipschitz continuity).
The transition function $f(\cdot)$ is $L_{f}$-Lipschitz-continuous, i.e.,
$\|f(z)-f(z^{\prime})\|_{2}{}\leq{}L_{f}d(z,z^{\prime}).$
@@
The individual policies $\pi$ present in any admissible policy profile $\bm{\pi}$ in $\Pi$ are $L_{\pi}$-Lipschitz-continuous, i.e.,
$\|\pi(s,\mu)-\pi(s^{\prime},\mu^{\prime})\|_{2}{}\leq{}L_{\pi}(\|s-s^{\prime}\|_{2}{}+{}W_{1}(\mu,\mu^{\prime}))$
for all $\pi\in\bm{\pi}\in\Pi$.
@@
The reward function $r(\cdot)$ is $L_{r}$-Lipschitz-continuous, i.e.,
$\|r(z)-r(z^{\prime})\|_{2}{}\leq{}L_{r}d(z,z^{\prime}).$
These assumptions are considered standard in model-based learning Jaksch et al. (2010); Chowdhury and Gopalan (2019); Sessa et al. (2022); Pásztor et al. (2023) and mild, as individual policies and rewards are typically designed such that they meet these smoothness requirements. For example, we use neural networks with Lipschitz-continuous activations to represent our policies (see Section D.3).
3.2. Examples of Safety Constraints
We can model multiple classes of safety constraints $h_{C}(\cdot){}\geq{}0$ that naturally appear in real-world applications such as vehicle repositioning, traffic flow, congestion control, and others.
Entropic safety. Entropic constraints can be used in multi-agent systems to prevent overcrowding by promoting spatial diversity and avoiding excessive clustering. Incorporating an entropic term in the decision-making process encourages the controller to distribute the agents evenly within the state space. This might be particularly useful in applications that include crowd behavior, such as operating a swarm of drones or a fleet of vehicles. In such scenarios, we define safety by imposing a threshold $C\geq 0$ on the differential entropy
$H(\mu){}:={}-\int_{\mathcal{S}}\log\mu(s)\mu(ds)$ (5)
of the mean-field distribution $\mu$, i.e.,
$h_{C}(\mu){}:={}H(\mu)-C.$
Distribution similarity. Another way to define safety is by preventing $\mu$ from diverging from a prior distribution $\nu_{0}$. The prior can be based on previous studies, expert opinions or regulatory requirements. We can use a penalty function that quantifies the allowed dissimilarity between the two distributions
$h_{C}(\mu;\nu_{0}){}:={}C-D(\mu,\nu_{0}),$
with $C\geq 0$ and where the distance function between probability measures $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{\geq 0}$ depends on the problem at hand.
We provide further examples of safety functions in Appendix B together with proofs that they satisfy Assumption 6.
3.3. Statistical Model and Safety Implications
The representative agent learns about unknown transitions by interacting with the environment. We take a model-based approach to achieve sample efficiency by sequentially updating and improving the transition model estimates based on the previously observed transitions. At the beginning of each episode $n$, the representative agent updates its model based on $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t{}+{}1})\}_{t=0}^{T-1}$ and $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$ is the set of observations in episode $i$ for $i=1,...,n-1$, i.e., up until the beginning of episode $n$. We estimate the mean $\bm{m}_{n-1}:\mathcal{Z}\rightarrow\mathcal{S}$ and covariance $\bm{\Sigma}_{n-1}:\mathcal{Z}\rightarrow\mathbb{R}^{p\times p}$ functions from the set of collected trajectories $\cup_{i=1}^{n-1}\mathcal{D}_{i}$, and denote model’s confidence with $\bm{\sigma}^{2}_{n-1}(z)=\text{diag}(\bm{\Sigma}_{n-1}(z))$. We assume that the statistical model is calibrated, meaning that at the beginning of every episode, the agent has high probability confidence bounds around unknown transitions. The following assumptions are consistent with Srinivas et al. (2010); Chowdhury and Gopalan (2019); Curi et al. (2020); Sessa et al. (2022); Pásztor et al. (2023) and other literature which aims to exclude extreme functionals from consideration.
@@ 3.3. Statistical Model and Safety Implications
Let $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ be the mean and covariance functions of the statistical model of $f$ conditioned on $n-1$ observed episodes. For the confidence function $\bm{\sigma}_{n-1}(\cdot)$, there exists a non-decreasing, strictly positive sequence $\{\beta_{n}\}_{n{}\geq{}0}$ such that for $\delta>0$ with probability at least $1-\delta$, we have jointly for all $n{}\geq{}1$ and $z\in\mathcal{Z}$ that $|f(z)-\bm{m}_{n-1}(z)|{}\leq{}\beta_{n-1}\bm{\sigma}_{n-1}(z)$ elementwise.
@@
The confidence function $\bm{\sigma}_{n}(\cdot)$ is $L_{\sigma}$-Lipschitz-continuous for all $n{}\geq{}0$, i.e., $\|\bm{\sigma}_{n}(z)-\bm{\sigma}_{n}(z^{\prime})\|_{2}{}\leq{}L_{\sigma}d(z,z^{\prime})$.
Since the true transition model is unknown in Equation 4c and Equation 4d, at the beginning of every episode $n$, the representative agent can only construct the confidence set of transitions $\mathcal{F}_{n-1}$ with $\bm{m}_{n-1}(\cdot)$ and $\bm{\sigma}_{n-1}(\cdot)$ estimated based on the observations up until the end of the previous episode $n-1$, i.e.,
$\mathcal{F}_{n-1}{}={}\Big{\{}\tilde{f}:\tilde{f}\text{ is calibrated w.r.t. }\bm{m}_{n-1}(\cdot)\text{ and }\bm{\sigma}_{n-1}(\cdot)\Big{\}}$ (6)
The crucial challenge in Safe Mean-Field Reinforcement Learning is that the representative agent can only select transitions $\tilde{f}\in\mathcal{F}_{n-1}$ at the beginning of the episode $n$ and use it instead of true transitions $f$ when solving Section 3.1 to find an optimal policy profile $\bm{\pi}_{n}^{*}$. The resulting mean-field distributions $\{\tilde{\mu}_{t}\}_{t=1}^{T}$ are then different from $\{\mu_{t}\}_{t=1}^{T}$ (i.e., the ones that correspond to the true transition model), and hence the constraint Equation 3 guarantees only the safety under the estimated transitions $\tilde{f}$, i.e., $h_{C}(\tilde{\mu}_{t}){}\geq{}0$. In contrast, the original environment constraint $h_{C}(\mu_{t}){}\geq{}0$ might be violated, resulting in unsafe mean-field distributions under true transitions $f$.
Next, we demonstrate how to modify the constraint Equation 4e for the optimization problem Section 3.1 when an estimated transition function $\tilde{f}$ is used from the confidence set $\mathcal{F}_{n-1}$ such that the mean-field distributions $\mu_{n,t}$ induced by the resulting policy profile $\bm{\pi}_{n}^{*}$ do not violate the original constraint under true transitions $f$. First, we require the following property for any safety function $h(\cdot)$.
@@
The following lemma shows that we can ensure safety under true transitions $f$ by having tighter constraints under any estimated transitions $\tilde{f}$ selected from $\mathcal{F}_{n-1}$.
@@
Given a fixed policy profile $\bm{\pi}_{n}$, a safety function $h(\cdot)$ satisfying Assumption 6 and a safety threshold $C\in\mathbb{R}$, we have in episode $n$ for all steps $t$
@@
For arbitrary $\tilde{\mu}_{n,t},\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ we have
$|h_{C}(\tilde{\mu}_{n,t}){}-{}h_{C}(\mu_{n,t})|{}={}|h(\tilde{\mu}_{n,t}){}-{}h(\mu_{n,t})|{}{}\leq{}{}L_{h}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t}){}{}\leq{}{}L_{h}C_{n,t},$
where the first equality follows from the definition of $h_{C}(\cdot)$, the first inequality follows from Assumption 6 and the second inequality comes from $C_{n,t}{}\geq{}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$.∎
Crucially, using Lemma 1 we can formulate a safety constraint for the optimization under estimated transitions $\tilde{f}$ that ensures that the constraint under true transitions $f$ is satisfied with high probability.
@@
For every episode $n$ and step $t$, $h_{C}(\tilde{\mu}_{n,t}){}\geq{}L_{h}C_{n,t}$ implies $h_{C}(\mu_{n,t}){}\geq{}0$ guaranteeing the safety of the original system.
@@
The corollary follows directly from Lemma 1 and the triangle inequality, which are used in the third and the second inequality, respectively
$L_{h}C_{n,t}$ ${}\leq{}h_{C}(\tilde{\mu}_{n,t})$
${}\leq{}|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|{}+{}h_{C}(\mu_{n,t})$
${}\leq{}L_{h}C_{n,t}{}+{}h_{C}(\mu_{n,t}).$
The claim is obtained by subtracting the positive constant $L_{h}C_{n,t}$ from both sides. ∎
@@
Then, $C_{n,t}$ for $t=1,\dots,T$ become parameters of the optimization problem (as defined in Section 4) that the representative agent faces at the beginning of episode $n$. However, choosing the appropriate values that comply with the condition $C_{n,t}{}\geq{}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ is not trivial since $\mu_{n,t}$ depends on unknown true transitions of the system. Note that computing $C_{n,0}$ at the initial step $t=0$ is not necessary because the inequality is always guaranteed due to the initialization $\tilde{\mu}_{n,0}=\mu_{n,0}$ for every episode $n$. In Appendix A, we demonstrate how to efficiently upper bound $W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and obtain $C_{n,t}$ using the Lipschitz constants of the system and the statistical model’s epistemic uncertainty. In particular, $C_{n,t}$ approaches zero, and $h_{C}(\tilde{\mu}_{n,t}){}\geq{}L_{h}C_{n,t}$ reduces to the constraint Equation 4e as the estimated confidence $\bm{\sigma}_{n-1}(\cdot)$ shrinks due to the increasing number of observations available to estimate true transitions.
Algorithm 1 Model-Based Learning Protocol in Safe-M3-UCRL
@@
3:     Compute $C_{n,t}$ for $t{}={}1,\ldots,T$ as described in Appendix A
4:     Optimize the objective in Section 4 over the admissible policy profiles $\Pi$ and hallucinated transitions Equation 8
5:     Execute the obtained policy profile $\bm{\pi}_{n}^{*}$ and collect the trajectories $\mathcal{D}_{n}=\{(z_{n,t},s_{n,t{}+{}1})\}_{t=0}^{T-1}$ from the representative agent
6:     Update the confidence set of transitions $\mathcal{F}_{n-1}$ with the collected data to obtain $\mathcal{F}_{n}$ for the next episode